solve the equation. Round to the nearest hundreth. check the rounded solution.
3(3t - 14 ) = -4

Answer by checkley75(3666)

3(3t-14)=-4
9t-42=-4
9t=-4+42
9t=38
t=38/9
t=38/9 answer.